### v4

# Rotation Curve Fitting Model

Abstract
One key piece of evidence for dark matter is the flat rotation curve problem: the disagreement between measured galactic rotation curves and their luminous mass. A novel solution to this problem is presented here. A model of relativistic frame effects on Doppler shifts due to the slightly curved frames of emitting galaxies with respect to the Milky Way frame is derived. This model predicts observed Doppler shifted spectra based only on the luminous matter estimates and one free model parameter. The rotation curve fitting model presented is tested on the SPARC sample of 175 galactic rotation profiles, also fitted by also fitted by dark matter models, and MOND via the Radial Acceleration Relation. For the SPARC sample, the new rotation curve fitting model gives an averaged reduced for 172 galaxies fitted, the isothermal dark matter model gives for 165 galaxies fitted, and the Radial Acceleration Relation (MOND precursor) gives for 175 galaxies fitted. The model presented here has a free parameter that is highly correlated with a ratio of photometric quantities. Implications of this model are discussed.
keywords:
gravitation < Physical Data and Processes – (cosmology:) dark matter – galaxies: fundamental parameters – galaxies: haloes – galaxies: bulges – galaxies: photometry1 Introduction
The flat-rotation curve problem is the divergence of two rotation velocities about the center of a spiral galaxy, inferred from different observations of light (Rubin et al., 1980; Bosma, 1981; van Albada et al., 1985): photometry and Doppler shifted spectra. Photometry gives estimates of the luminous mass, which when interpreted classically by the Poisson equation gives the expected Keplerian orbital velocities which decline beyond the stars. Doppler shifted spectra, however, give the “flat-rotation curve” velocities, which remain essentially constant far past the stars. The divergence of these two velocities has been primary evidence for dark matter theories (Fig. 1).
Dark matter is hypothesized to be massive particles which are electromagnetically neutral, and have a very low interaction probability with baryonic matter, but otherwise obey classical gravity. Though direct detection experiments continue, in the absence of a definitive observation of such particles (Cebrián, 2022), the phenomenological details of dark matter problems remain interesting. S. McGaugh notes a degeneracy between the luminous and dark matter components in a galaxy, when he asks “Why is the luminous tail wagging the dark matter dog, if dark matter dominates dynamics?” (McGaugh, 1999). This references the curious fact that knowledge of the stellar disk completely determines the spherical dark matter halo, even though dark matter is postulated to dominate dynamics (McGaugh, 2004).
Another interesting trend is the Universal Rotation Curve (URC), in which a spectrum of 1,100 rotation curves (RCs) inflect about the presumed rotation curve of the Milky Way (Rubin et al., 1978; Persic et al., 1996; Salucci et al., 2007). In the URC spectrum, galaxy RCs are separated with respect to the luminous mass of the Milky Way, where the RCs of galaxies larger than the Milky Way inflect downwards towards “flat”, and RCs of galaxies smaller than the Milky Way inflect upwards towards “flat”. The Milky Way has been postulated to be at the inflection point where truly flat RCs lie, though a definitive rotation curve past our position at approximately kpc has been difficult to obtain until the current era of observations from Gaia DR3 (Jiao, Yongjun et al., 2023). We interpret this positioning of the Milky Way as evidence for frame dependent effects in this problem.
Dark matter theory accounts for the curious URC phenomenology with the phrasing that galaxies smaller than the Milky Way are “dark matter dominated”, and those larger than the Milky Way require only “minimal dark matter halos”. Dark matter particles are designed to obey classical gravity, and so this logical inconsistency requires fine-tuning of extra free-parameters, reducing the predictivity of the models (McGaugh, 2021). In classical gravity, mass accretion rates are directly proportional to the initial mass function (Maschberger et al., 2014).
In this paper we interpret the URC phenomenology, instead, as due to relative frame effects from our Milky Way, which immediately explains our apparent position in the middle of a spectrum of 1,100 rotation curves. This interpretation of the problem removes extraneous free parameters, and does not modify classical gravity theory. Modified Newtonian Dynamics (MOND) (Milgrom, 1983), the leading alternative to dark matter, is similar to what we present in so far as MOND states that the luminous mass is the only mass, but differs from our model as it modifies classical gravitational physics to obtain rotation curve velocities. We will instead interpret excess rotation velocities, above the estimates from photometry, as due to an incomplete interpretation of Doppler shifted spectra, not true orbital speeds.
MOND is instructive in this context because it successfully fits a diverse distribution of galaxy rotation curves (RCs), by modifying the standard gravitational acceleration scale on the length scales of galaxies (McGaugh, 2014). Previous attempts to extend MOND into the relativistic regime have given rise to new theories of gravity (Bekenstein, 2004; Cooperstock & Tieu, 2007; Famaey & McGaugh, 2012), but what we propose is simpler, a transition of the concept of a changing acceleration to the relativistic regime of a changing relative curvature. It is remarkable that MOND’s free parameter is essentially constant across a wide range of galaxy morphologies (Lelli et al., 2016), which we interpret as a characterization of the role of the Milky Way in this problem. However, there are some notable exceptions where MOND fits of well-studied galaxies indicate distances which are in conflict with standard candles, and this problem is removed in our formulation of the flat-rotation curve problem.
The new rotation curve fitting model (RCFM) presented here fits the same galaxy data as MOND and dark matter models, at reported distances, but with fewer free parameters and no modification of physics. The RCFM model is a one-parameter fit to the rotation curve data, whose only inputs are estimates of the luminous mass of the galaxy emitting the photons and of galaxy receiving those photons. Recent observations of the Milky Way from the Gaia DR3 (Jiao, Yongjun et al., 2023) indicate a Keplerian decline in rotation velocities from kpc to kpc, consistent with the model paradigm presented here.
The new rotation curve fitting model (RCFM) is derived in Section 2. Historically, gravitational redshifts considered in this problem were obviated by Galilean subtraction of redshifts at the large limit of the data of the emitter galaxy from the receiver galaxy (Misner et al., 1973). We instead employ a relativistic framing, in which field frames are parametrized by the Schwarzschild gravitational redshift terms, and then we Lorentz map between between the emitter and receiver galaxies one-to-one in radius (See Section 2.4). In Section 3 we describe the SPARC database (Lelli et al., 2016) of 175 well known galaxy rotation curves and associated luminous mass models, and the Milky Way baryon models used in this work. In Section 4 we compare RCFM fit results to those from dark matter models and the Radial Acceleration Relation (RAR) (McGaugh, 2014; McGaugh et al., 2016; Lelli et al., 2016; Li et al., 2018) for the SPARC galaxy database (RAR has a simpler functional form and is a possible precursor for MOND based on the same physical paradigm (Dutton et al., 2019)). In this section we also explore a correlation between the RCFM free parameter and a ratio of photometric observables (luminosity and half light radius of the galaxy). In Section 5 we conclude with discussion of upcoming observations which can further constrain or falsify this model.
2 Rotation Curve Fitting Models
2.1 Two observations of light
The observables in the flat-rotation curve problem are the Doppler shifted spectra and photometry.
Doppler shifted spectra is interpreted as RC velocities by the Lorentz boost
for the characteristic lab frequency, the observed shifted frequency, and , for the vacuum light speed. We emphasize this form of the Lorentz boost is applicable to a flat, inertial spacetime.
The second observable, photometry, gives a measure of total light which is then interpreted as mass from population synthesis models (PSMs) which produce mass-to-light ratios and hence models of the baryon distribution in a given galaxy. PSM rely upon a complex suite of assumptions regarding galaxy evolution, metallicities and initial mass functions (Schombert et al., 2018; Bell & de Jong, 2001), and are under-constrained due to the dark matter problem (Conroy et al., 2009; Dutton et al., 2005), producing an error budget of (Lelli et al., 2016).
Mass models give velocities by parametrizing the integration of the Newtonian gravitational potential by
for the field point, the reference point, and the acceleration due to the gravitational field. The classical boundary condition is that the potential goes to zero at , hence the integration effectively starts from large limit of the data, and integrates in to the small . This potential then solves the Poisson equation
where is the mass density distribution and is Newton’s constant. The gradient of this potential then gives the circular orbital velocities (Eq. 6) by the central force relation
The baryonic components of a galaxy include the stellar disk, stellar bulge, and a gas halo, which are summed in quadrature to reflect a sum of gradients
Mass-to-light ratios and are free parameters in most rotation curve fits due to the under-constrained nature of luminous mass modeling. Gas fractions are calculated from a different observational technique (Casertano, 1983), and do not require mass-to-light ratio modeling.
2.2 Dark matter rotation curve fitting formula
The dark matter rotation curve (RC) formula is formulated in the same sense as Eq. 5, as a sum of gradients in the potential
for terms in the Keplerian velocity predictions (Eq. 5), the velocities attributed to dark matter, and the model predictions which are fitted to the RC velocities reported from observations of Doppler shifted spectra (Eq. 1). All velocities are assumed to be those of test particles in circular orbits, in the plane of the stellar disk, about the rotation axis of the galaxy at .
2.3 Radial Acceleration Relation
The Radial Acceleration Relation (RAR)(McGaugh et al., 2016) has been considered as the phenomenological basis for MOND (Dutton et al., 2019; Desmond et al., 2023), as it has a simpler functional form and less scatter than MOND but the same physical paradigm. For this reason, we will compare RCFM fits to RAR and dark matter fits to the SPARC sample. The RAR approaches rotation curves from the MONDian paradigm of a changing law of inertia, but has also been shown to be consistent with dark matter fits on a sample of simulated galaxies (Keller & Wadsley, 2017).
2.4 New rotation curve fitting model (RCFM)
The physical paradigm for the new rotation curve fitting model (RCFM) is: the luminous mass is the only mass, the velocity due to the baryons is the only velocity, and excesses in Doppler shifted spectra represented by the “flat” rotation velocities are due to relative frame effects from the Milky Way galaxy. For clarity in what follows, we now drop explicit functional dependence on and set constants like Newton’s and the speed of light to 1. All terms can be assumed to be evaluated one-to-one in radius between the galaxy being observed and the Milky Way, with the exception of the model’s free parameter which is single valued for each galaxy fitted.
We construct a new heuristic rotation curve fitting formula by replacing the dark matter contribution to the rotation curve formula in Eq. 6 with the convolution of two Lorentz-type maps and which represent the frame-dependent effects due to the Milky Way;
This prescription is an implicit assumption that contributions to Doppler shifted spectra from translation ( in Eq. 5) and curvature () are roughly separable (Jackson, 1999) as represented by Eq. 7.
Terms in are a measure of relative curvature
The RCFM galaxy maps and (Eq. 7) are constructed from the tetrad formalism of general relativity (Tecchiolli, 2019), where an inertial frame field metric (Minkowski ) can be attached to any point on a curved manifold. The frame fields are related to the curved spacetime metric in the following way
In the frame field formalism “when going from one local inertial frame at a given point to another at the same point, the fields transform with respect to a Lorentz transformation” (Jetzer, 2017). Our goal then, is to use Lorentz-type transformations to map target galaxies onto the Milky Way, one-to-one in radius. To do this, we use the weak-field Schwarzschild time metric coefficients to characterize the frame fields, and hence the effects on photon frequency. The weak-field timelike Schwarzschild field frames are defined by
Terms in are the Newtonian gravitational potential (Wald, 1984). Classically, the of galaxies are integrated from the large limit of the data with a boundary value of , into a negative maxima at the small limit (Eq. 2). This is an implicit assumption of a globally flat embedding space, which ensures that the potential goes to zero at a distance of infinity from the mass distribution.
However, recent observations (Pomarède et al., 2020) have shown that on the relevant length scales of galaxies and groups of galaxies we have not recovered flat spacetimes. In fact, the external environments of galaxies and flowlines of groups of galaxies are exceedingly complex and diffusely populated with matter, demonstrating structure with sources and sinks. In addition, the value of the vacuum energy remains an outstanding physics problem (Zel’dovich, 1968).
In the absence of definite knowledge of the values of at the large limit of individual galaxies,
and because all RCFM terms are ratios of terms in for the emitter and receiver galaxies, we instead want to compare galaxies from a position where we know they have the same curvature, namely at their centers. To inform the integration of the gravitational potentials, we note Wolfgang Rindler’s statement that “the center of each galaxy provides a basic local standard of nonacceleration … so then can be treated like a local inertial frame relative to its own center”(Rindler, 2013).
From Rindler’s insight, we propose that to compare galaxies as inertial frames by Lorentz-type maps we must integrate from their centers out to the limit of the data.
In practice, this means that galaxy potentials are summed from at out to an essentially constant value
at the large limit of the data. This produces positive definite gravitational potentials,
which are then subtracted from in the Schwarzschild clock term .
Gravitational
potentials calculated in this way still obey Poisson’s equation Eq. 3 and the central force law Eq. 4, but allow for a non-flat embedding space with no loss of generality.
The RCFM mapping is accomplished in two steps. First, the curved 2-frame mapping maps the entire emitter galaxy onto the Milky Way one-to-one in radius,
for a rapidity defined by
where represents the emitter galaxy and the Milky Way. A cartoon of the transformation is represented in Fig. 2.
The second transformation takes this curved 2-frame (Eq. 12) into the flat field frames where observations are made and where the metric is Minkowski , as
for a rapidity angle defined by
for a convolution of the curved frame (Eq. 12) with the flat frame
Terms in are the Keplerian velocities from luminous mass as in Eq. 5. A cartoon of the transformation is represented in Fig. 3 .
That the Keplerian velocities are our best estimate of flatness is evidenced by the fact that dark matter is not required to reproduce the RC of our Solar System.
Previously, Rindler extended the transforms of special relativity to accelerated frames on a flat background (Boyer, 2011). Viewed as Rindler’s accelerated coordinates, and respectively can be seen to be timelike and spacelike transformations (Misner et al., 1973; Wald, 1984; Rindler, 2013).
2.5 Geometric simplifications
Population synthesis models commonly assume spherical symmetry for the stellar bulge and gas halo, but axial symmetry for the stellar disk (Schwarzschild, 1954; Freeman, 1970). However, it is a common calculational tool to use spherical symmetry for the entire integrated potential of the luminous mass distribution (Bekenstein, 2004; McGaugh, 2008; Mistele et al., 2022) because numerical integration of the thin disk is computationally intensive, requiring assumptions of under-constrained boundary conditions and relevant physical scales, which therefore add extra free parameters to the problem (Huré & Hersant, 2011). We use the spherically symmetric Schwarzschild metric here as proof of concept, as it is the simplest representation which captures the relevant physics of the problem without excess computations.
For a given mass, the gravitational potentials calculated in a spherically symmetric geometry converge to those for an exponential disk geometry at lengths greater than one-third of the exponential scale-length , ie. (Chatterjee, 1987). Since this is the region where dark matter effects becomes important (van Albada et al., 1985), this calculational technique captures the relevant physics. However, in the region where the RCFM integrations begin galaxy comparisons ( to ) the spherical assumption overestimates the potential by a factor of , which can be seen in RCFM disk mass-to-light ratio results. Implementation of a thin disk geometry would resolve this problem (de Vaucouleurs, 1959),(Freeman, 1970), though at significant computational expense.
3 Data
3.1 SPARC galaxies
We fit the Spitzer Photometry and Accurate Rotation Curves (SPARC) dataset of 175 nearby galaxies with extended RC data from atomic hydrogen (HI) and H- (Lelli et al., 2016). HI provides the most reliable RCs because it is dynamically cold, traces circular orbits, and can be observed several effective radii past the stellar disk. This sample of rotationally supported galaxies spans the widest range of masses and morphologies presently available.
These galaxies are accompanied by Keplerian velocities which represent the baryonic mass model from PSM, based on Spitzer Photometry in the near infrared at 3.6. Near infrared is currently believed to be the best tracer of stellar mass in population synthesis models (PSM) (Schombert et al., 2018), as at this wavelength, mass-to-light () ratios are believed to be almost constant and independent of star formation history (Bell & de Jong, 2001; Schombert et al., 2018). The SPARC database reports mass-to-light ratios of in units of at 3.6. Gas fractions are calculated from surface density profiles of HI with the formalism given in (Casertano, 1983) and scaled by a factor 1.33 to account for cosmological helium abundances. Contributions from molecular gas are ignored because CO data are not available for most SPARC galaxies. Error on these velocities is estimated at (Lelli et al., 2016). The SPARC database can be found at http://astroweb.cwru.edu/SPARC/.
3.2 Milky Way Luminous Mass Models
The RCFM, as currently formulated, requires a static choice of a luminous mass distribution for the Milky Way. Determining the luminous mass profile of the Milky Way (MW) is an under-constrained problem, due to our observing position from within the galactic disk (Fich & Tremaine, 1991). In this paper, we compare two different Milky Way baryon models to the SPARC sample; one from McGaugh (McGaugh, 2008) and one from Sofue (Sofue, 2013).
The two MW models differ markedly in the inner kpc; as the McGaugh MW has a triaxial bar-bulge and the Sofue MW has a de Vaucouleurs bulge. The Sofue MW model is used from to kpc, and extended with the Xue MW model from to kpc (Xue et al., 2008). The Sofue and Xue MW models come from the same data and a dark matter model is used for the extended rotation curve. The McGaugh Milky Way model covers from to kpc, comes from data, and the MOND model is used for the extended rotation curve.
The recent ESA Gaia data release (DR3), Jiao, Yongjun et al. (2023) demonstrates a Keplerian decline of the rotation curve of the Milky Way in the range of to kpc, consistent with this RCFM heuristic paradigm. The ESA Gaia mission, taking measurements from the second Lagrange point, has revolutionized the science of the MW with unprecedented detail, statistical accuracy, and a drastic reduction in systematic uncertainties. Since this is the most recent and best data for the MW, we replace the model Milky Way model velocities in the range from to kpc with those from Jiao, Yongjun et al. (2023). We then shift the two remaining sides of the given MW model by a constant amount to match the endpoints of the Jiao rotation curve. See Table 1 for summary details.
| Author | scale length | Model for the | Range |
| bulge/bar | extended RC | ||
| Sofue-Xue | 0.5 kpc | NFW (Navarro et al., 1996) | [0,60] kpc |
| Sofue (2013); Xue et al. (2008) | bulge | dark matter | |
| McGaugh | 2.0 kpc | MOND | [0,150] kpc |
| McGaugh (2008) | bar | (Milgrom, 1983) |
4 Analysis and Results
4.1 Fitting procedure
To fit galaxies in the SPARC sample with the RCFM model, the fitting procedure is as follows. First, a Milky Way baryon model is selected (Sec. 3.2) and the model data in is read in for a series of measurements in radii. The galactic gravitational potential is calculated by numerically integratingas (see Sec. 2.4),
Once the MW potential is calculated, it remains static for the rest of the fitting procedure.
The data for the galaxies being observed include several pieces of information: the RC velocities from Doppler shifted spectra, the uncertainty on that measurement , and the components of the luminous mass interpreted as orbital velocities, , , and . To calculate the baryonic potential for the galaxy in question, is first computed as per Eq. 5, for and . The associated with that galaxy is then computed as in Eq. 16.
After the for the galaxy being studied and the Milky Way have been calculated, the components must be compared at matching values of . To match radii, is interpolated to produce values at the radii reported in the measured RC data of the galaxy being observed. Any point with a radius larger than the largest radius in the Milky Way model is discarded.
The RCFM prediction is assembled as in Eq. 7, to give a predicted which is compared to the RC data . The equations outlined in Section 2 contain free parameters that must be determined for each galaxy fit: , , and . The model’s free parameter starts from an initial value of , and mass-to-light parameters from an initial value of , and then all parameters are allowed to vary freely and to be determined by minimization of the . The scipy.optimize.curve_fit utility in Python is used to perform this minimization.
4.2 Evaluating goodness-of-fits
There are two metrics by which different rotation curve fitting models are compared, the resulting reduced values and the mass-to-light ratios. Since error estimates on RC velocities have not been standardized across the field (de Blok et al., 2008a; Gentile et al., 2011), values can only be compared when fitted to the same RC data. In Table LABEL:table:M2Light we compare values from three models: dark matter halo, Radial-Acceleration-Relation (RAR) (McGaugh et al., 2016; Lelli et al., 2017; Li et al., 2018) and RCFM, fitted to the SPARC database of galaxy RCs (Lelli et al., 2016) . RCFM values are remarkably low, providing confidence in the faithfulness of the model to galaxy RC data.
The average mass-to-light ratios from RCFM fits yield bulge mass-to-light ratios which are within the error estimates on population synthesis models (PSM). The RCFM disk mass-to-light ratios are approximately a factor of two larger than PSM, consistent with the artifact introduced from the spherical integration technique, see Sec. 2.5 for a full explanation. This artifact can be removed by including an exponential disk integration at the cost of additional computation.
4.3 Comparing Milky Way models
RCFM fits, as currently formulated, require a choice of a static Milky Way baryon distribution. In Table 9 and LABEL:TSet we compare the RCFM fit results for the two different Milky Way model assumptions used in this paper. As can be seen in the table, for the 36 most reliable galaxies in the SPARC sample the two Milky Way models are equivalent (see Sec. 4.6). However, for the whole SPARC sample of 175 galaxies the RCFM fits using the Sofue-Xue-Jiao Milky Way performs better with an average , versus those from McGaugh-Jiao Milky Way . Based on this, values in Table LABEL:table:M2Light and figures are from RCFM fits assuming the Sofue-Xue-Jiao Milky Way.
The residuals of fits to the SPARC sample are also used to compare the two different MW models (see Sec. 3.2). Histograms of residuals normalized by the error in velocity observations are shown in Fig. 4. In all cases, residuals of model fits to observed velocity data followed a narrow distribution centered at zero with a range of standard deviations in error, albeit with heavy tail features. The behavior of the residuals did not vary greatly between MW models, suggesting that the fitting parameters in our model are robust with respect to differing MW model assumptions at this present level of analysis. With a finer grained analysis and a larger sample, the degeneracy between MW models may be resolved.
A Gaussian fit on residuals from fits using the McGaugh-Jiao MW model gave a mean of , standard deviation of in units of standard deviations in velocity error; whereas a Gaussian fit to the residuals from fits using the Sofue-Xue-Jiao MW model gave a mean of , standard deviation of in those same units. The residuals are shown in Fig. 4. The small values associated with these quantities in both cases provide confidence that our fits match data closely.
However, the Gaussian fits did not quite capture the full peak and heavy tails in the residuals. This suggests that there may be non-Gaussian error in the observations of galaxy velocities. To address this, the residuals were also fit to an exponential function shown in Fig. 5. The exponential function captured both the peak and the heavy-tailed behavior of the distributions more faithfully.
4.4 Individual Galaxy Results
In this section we will compare fit results from the RCFM, RAR and dark matter models for well-studied galaxies. All reported in what follows are in units of solar mass per solar luminosity .
4.4.1 NGC 2841
NGC 2841 is a star dominated, flocculent spiral galaxy, which historically has been regarded as a problematic case for MOND ((Gentile et al., 2011)). RAR finds a good fit for this galaxy after adjusting the Cepheid distance of Mpc by , to Mpc. Cepheid variable stars are the most accurate distance indicators currently available. The RCFM fit is shown in Fig. 6 at the Cepheid based distance, and fit results are compared to dark matter and RAR in Table 2.
| Model | distance(Mpc) | ||||
| Dark Matter Iso | 0.60 | 0.66 | 1.58 | 14.1 | |
| RAR | 0.81 | 0.93 | 1.52 | 15.5 | |
| RCFM | 0.91 | 1.10 | 1.36 | 14.1 |
4.4.2 IC 2574
IC 2574 is a dwarf spiral galaxy, dominated by gas with no central stellar bulge. RAR finds a good fit to this galaxy but adjusts the distance and inclination by to . This galaxy is problematic for dark matter model fits, as the model overestimates the inner RC out to 10 kpc (Navarro et al., 2017). RCFM fits this galaxy successfully, at the reported tip of the red giant branch distance of Mpc, Fig.7, model comparisons are in Table 3.
| Model | ||||
| Dark Matter Iso | 0.77 | – | 2.51 | |
| RAR | 0.07 | – | 1.44 | |
| RCFM | 1.10 | – | 2.27 |
4.4.3 NGC 3198
NGC 3198 is a barred spiral in Ursa Major, which has been well studied. It has been considered a problem galaxy for MOND, when the distance is a free parameter in fits (Gentile et al., 2011), as the preferred MOND distance is 2 different from that reported from Cepheids. However, in the SPARC database the RAR fit to this galaxy at the Cepheid distance reproduces the RC well. This galaxy has no bulge component, as indicated in Table 4. See Fig. 8 for the RCFM fit.
| Model | distance (MPC) | ||||
| Dark Matter Iso | 0.52 | – | 1.31 | 13.7 | |
| RAR | 0.77 | – | 2.06 | 13.8 | |
| RCFM | 0.88 | – | 1.72 | 13.8 |
4.4.4 NGC 7814 and NGC 891
The spiral galaxies NGC 7814 and NGC 891 present an interesting challenge to dark matter and MOND models. Both galaxies are presented edge-on on the sky, and both have essentially identical RCs, but are extreme opposites in morphologies. NGC 7814 is a bulge dominated galaxy and NGC 891 is almost entirely a disk galaxy. Fraternali et al. (2011) asks “why are these RCs so identical if their dark matter halos are necessarily different to accommodate the differences in the luminous mass?”
In the RCFM paradigm the two RCs are very similar in magnitude, but differ in their inflections at large radii; NGC 7814 inflects up, whereas NGC 891 inflects down. RCFM fits to these galaxies are successful, see Fig.9 and Tables 5 and 6. We report here fits to the (Fraternali et al., 2011) RC data included in the SPARC database.
| Model | ||||
| Dark Matter Iso | 0.68 | 0.71 | 0.39 | |
| RAR | 1.17 | 0.52 | 1.334 | |
| RCFM | 0.38 | 0.68 | 0.63 |
| Model | ||||
| Dark Matter Iso | 0.77 | 1.63 | 1.30 | |
| RAR | 0.32 | — | 25.16 | |
| RCFM | 0.66 | — | 1.89 |
4.4.5 NGC 5055
In this section we compare two different sets of NGC 5055 RC data from HI, and their resulting dark matter, RAR and RCFM fits. The first RC is from (Battaglia et al., 2006) as reported in SPARC. Their luminous model has no stellar bulge. The RC data and luminous mass model are reported at the reliable distance of Mpc, from the Tip of The Red Giant Branch method.
| Model | ||||
| Dark matter - Isothermal Halo | 0.26 | - | 6.19 | |
| RAR | 0.56 | - | 7.42 | |
| RCFM | 0.62 | - | 6.29 |
The second RC for NGC 5055 is from the THINGS (de Blok et al., 2008a) at a distance of Mpc from the Hubble Flow, including a stellar bulge component. RCFM fits to both sets of RC data can be seen in Fig.10, at reported distances, using the reported luminous mass models. This galaxy serves as an example of the under-constrained nature of galaxy mass modeling (Conroy et al., 2009). Details of the fits are given in Table 10.
| Model | ||||
| Dark Matter Iso | 0.79 | 0.11 | 8.13 | |
| Dark Matter NFW | 0.79 | 0.11 | 10.31 | |
| RAR | 0.43 | 0.46 | 2.63 | |
| RCFM | 0.57 | 0.23 | 1.30 |
4.5 Galaxies for which model fits fail
Quality flags are assigned in the SPARC galaxy database using the following schema; are galaxies with high-quality rotation curves; are galaxies with minor asymmetries and/or rotation curves of lower quality; Q = 3 for galaxies with major asymmetries, strong non-circular motions, and/or offsets between HI and stellar distributions, so that galaxies are not suited to detailed dynamical studies. Of the three models compared in this paper against the 175 SPARC galaxies; RAR fits all galaxies, the RCFM fits 172 of 175 galaxies (fits fail for UGC06628 (a galaxy), and F561-1, UGC04305 ( galaxies)), and the isothermal dark matter model fits 165 of 175 galaxies (fits fail for galaxies D512-2, UGC00634, NGC6789, UGC00891, UGC02023, UGC05999, UGC07232 and UGC09992, and galaxies F567-2 and F574-2).
4.6 Free Parameter Correlation
To test a physical interpretation of the model’s free parameter with respect to a ratio of total luminosity by the half-light radius (effective radius), we select a subset of galaxies from the SPARC dataset by the following criteria:
1.
Select galaxies with the most accurate and reliable distance estimates (tip of the red giant branch and Cepheid variable stars), rejecting all other galaxies.
2.
Select galaxies with inclinations on the sky in the range of to , rejecting galaxies with an inclination greater than as difficult to assign a true surface brightness profile, and those at inclinations less than as being difficult to report line of sight Doppler shifts accurately.
3.
Exclude galaxies with quality factor as not suited to dynamical studies due to asymmetries, non-circular motions, and/or offsets between stars and gas, using assigned SPARC factors as in (Lelli et al., 2016).
By this process, a training set of galaxies are selected, as reported in Table 9. We then plot the subset’s values versus the ratio of , and fit the distribution with a power law as shown in Fig. 11. The total luminosity and effective radii numbers used are as reported in the SPARC database. Luminosity measurements are taken in the wavelength of , assuming a solar absolute magnitude of 3.24 at (Oh et al., 2008) in units of solar . As can be seen in in Fig. 11 the subset appears to be highly correlated to the ratio of photometric quantities, though slightly different functionally for the choice of Milky Way.
| Model | ||||
| RAR | 0.46 | 0.74 | 3.89 | |
| RCFM (MW: Sofue-Xue-Jiao) | 1.04 | 0.87 | 1.60 | |
| RCFM (MW: McGaugh-Jiao) | 1.00 | 0.87 | 1.65 | |
| Dark Matter | 0.50 | 0.66 | 1.59 |
A subset of galaxies selected based on best distance estimates and highest quality rotation curve data. Reduced values do not reflect galaxies which fail to converge on a fit; for which, the RCFM and RAR have none, and the isothermal dark matter model has two (NGC6789, UGC07232). The average stellar mass-to-light ratios are presented in units of solar mass per solar luminosity . The RCFM spherical simplification used in RCFM fits for computational and formulaic simplicity, can be seen to yield a disk mass-to-light ratio which is approximately a factor of larger than those from RAR and Dark Matter fits. This is a well known artifact of the spherical assumption, and could be removed with a full disk geometry at the cost of additional computations and free-parameters, see Sec. 2.5 for more information.
| MW: | MW: | ||||
| Galaxy | McGaugh-Jiao | Sofue-Xue-Jiao | Distance | Distance | Inc |
| Name | (Mpc) | Method 111 Distance method: 2 = tip of the red giant branch, 3 = Cepheids. Distance and inclination information from (Lelli et al., 2016). | (deg) | ||
| CamB | 0.52 | 0.36 | 3.36 | 2 | 65 |
| D564-8 | 0.03 | 0.07 | 8.79 | 2 | 63 |
| D631-7 | 0.52 | 0.36 | 7.72 | 2 | 59 |
| DDO154 | 8.65 | 8.66 | 4.04 | 2 | 64 |
| DDO168 | 4.17 | 4.12 | 4.25 | 2 | 63 |
| ESO444-G084 | 1.23 | 0.63 | 4.83 | 2 | 32 |
| IC2574 | 1.88 | 2.05 | 3.91 | 2 | 75 |
| NGC0024 | 0.82 | 0.80 | 7.3 | 2 | 64 |
| NGC0055 | 2.24 | 2.40 | 2.11 | 2 | 77 |
| NGC0247 | 2.26 | 2.24 | 3.7 | 2 | 74 |
| NGC0300 | 0.41 | 0.41 | 2.08 | 2 | 42 |
| NGC2403 | 12.05 | 11.52 | 3.16 | 2 | 63 |
| NGC2683 | 0.95 | 0.97 | 9.81 | 2 | 80 |
| NGC2841 | 1.51 | 1.35 | 14.1 | 3 | 76 |
| NGC2915 | 0.62 | 0.58 | 4.06 | 2 | 56 |
| NGC2976 | 0.45 | 0.46 | 3.58 | 2 | 61 |
| NGC3109 | 0.30 | 0.27 | 1.33 | 2 | 70 |
| NGC3198 | 1.42 | 1.50 | 13.8 | 3 | 73 |
| NGC3741 | 0.76 | 0.63 | 3.21 | 2 | 70 |
| NGC4214 | 1.55 | 1.32 | 2.87 | 2 | 15 |
| NGC5055 | 6.29 | 6.10 | 9.9 | 2 | 55 |
| NGC6503 | 1.14 | 1.14 | 6.26 | 2 | 74 |
| NGC6789 | 0.86 | 0.94 | 3.52 | 2 | 43 |
| NGC6946 | 2.08 | 1.98 | 5.52 | 2 | 38 |
| NGC7331 | 1.08 | 1.09 | 14.7 | 3 | 75 |
| NGC7793 | 0.72 | 0.72 | 3.61 | 2 | 47 |
| UGC04483 | 0.38 | 0.44 | 3.34 | 2 | 58 |
| UGC07232 | 1.07 | 1.02 | 2.83 | 2 | 59 |
| UGC07524 | 1.40 | 1.41 | 4.74 | 2 | 46 |
| UGC07559 | 0.17 | 0.20 | 4.97 | 2 | 61 |
| UGC07577 | 0.06 | 0.06 | 2.59 | 2 | 63 |
| UGC07866 | 0.08 | 0.07 | 4.57 | 2 | 44 |
| UGC08490 | 0.15 | 0.13 | 4.65 | 2 | 50 |
| UGC08837 | 0.62 | 0.65 | 7.21 | 2 | 80 |
| UGCA442 | 0.74 | 0.73 | 4.35 | 2 | 64 |
| UGCA444 | 0.10 | 0.10 | 0.98 | 2 | 78 |
5 Conclusions
At this time, there are many discrepancies between cosmology theory and observations which are resolved by dark matter models (Sanders, 2010; Tully et al., 2014; Naidu et al., 2022). In this paper we address only one such discrepancy, the flat-rotation curve problem of spiral galaxies. This choice is due to the clear symmetry presented in the rotation curve data. Other dark matter problems require cosmological model assumptions, beyond the scope of the current paper.
Here we reinterpret the flat-rotation curve problem not as a problem of missing mass but rather, at the level of the data, as a problem of relative frame effects due to our home galaxy. We predict that the flat-rotation curve problem is an artifact of misinterpretation of Doppler shifted spectra from external galaxies, removing the need for dark matter halos on galaxies. This framing places primary importance on the role of the Milky Way, and compactly explains both why the Milky Way sits roughly at the inflection point in the Universal Rotation Curve spectrum of 1,100 galaxies (Persic et al., 1996) and why MOND is both successful and limited.
The rotation curve fitting model (RCFM) presented here reproduces the fitting successes of MOND, RAR (Radial Acceleration Relation) and dark matter models on a sample of 175 well studied galaxy rotation curves (Lelli et al., 2016), but does not modify classical physics. Independent of the interpretation of the various quantities appearing in the formulae presented in this paper, this is a one parameter fit to the data. What is more, the RCFM free parameter is highly correlated with a ratio of observable photometric parameters, which makes this the first direct constraint to luminous mass modeling of galaxies from Doppler shifted spectra.
We emphasize that the only input to the RCFM is the luminous mass. In contrast to previous investigations of gravitational redshifts in this context, which used Galilean subtraction (Misner et al., 1973), this approach sets the frame comparison from the center of the galaxies, using Lorentz-type boosts between galaxy frames in radius. In this way, using gravitational redshifts to define frames, we can explain the discrepancy between the “flat-rotation” curves and the Keplerian luminous mass predictions with only one free parameter. Our results are presented in this paper in comparison to dark matter and RAR fit results for the same galaxy data, from the SPARC database of 175 rotation curves (Lelli et al., 2016). RAR has a simpler functional form than MOND, though the same physics paradigm. The recent Gaia DR3 data release Jiao, Yongjun et al. (2023) has shown that the rotation curve of the Milky Way demonstrates a Keplerian decline from - kpc, consistent with the RCFM model paradigm presented here. The upcoming Large Survey of Space and Time at the Vera C. Rubin Observatory (Željko Ivezić et al., 2019) can falsify this model if the Milky Way rotation curve beyond kpc is found to require a dark matter halo.
6 Acknowledgments
This work is dedicated to Emmett Till. We acknowledge and express gratitude to the first nations peoples on whose ceded and unceded lands this was written; including but not limited to the the Coast Salish bands of the Puget Sound, the Cheyenne and Arapaho Tribes and Ute Tribes of Colorado, the Algonquian and Iroquoian Peoples of Massachusetts and New York, and the Navajo and Pueblo Nations of New Mexico. The authors would like to thank V. P. Nair, T. Boyer, M. Kaku, I. Chavel, R. Walterbos, N. P. Vogt, S. McGaugh, A. Klypin, T. Quinn, S. Tuttle, M. Juric, R. Rivera, N. Oblath, J. Formaggio, J. Conrad, P. Fisher, Y. Sofue, C. Mihos, and M. Merrifield.
| MW:SXJ | Dark Matter | RCFM | RAR | |||||||
| Galaxy | Q | |||||||||
| Averages: | 1.90 | 0.51 | 0.11 | 2.39 | 1.13 | 0.78 | 4.22 | 0.64 | 0.73 | – |
| CamB | 2.99 | 0.34 | 0.00 | 0.23 | 7.91E-05 | … | 5.76 | 0.34 | … | 2.00 |
| D512-2 | NAN | … | … | 0.21 | 1.48 | … | 0.37 | 0.48 | … | 2.00 |
| D564-8 | 0.26 | 0.49 | 0.00 | 0.11 | 1.21 | … | 3.16 | 0.40 | … | 2.00 |
| D631-7 | 0.85 | 0.40 | 0.00 | 0.30 | 0.26 | … | 15.87 | 0.20 | … | 1.00 |
| DDO064 | 0.48 | 0.51 | 0.00 | 0.45 | 1.58 | … | 0.33 | 0.48 | … | 1.00 |
| DDO154 | 3.23 | 0.41 | 0.00 | 10.58 | 1.18 | … | 3.48 | 0.19 | … | 2.00 |
| DDO161 | 0.36 | 0.50 | 0.00 | 0.62 | 0.97 | … | 1.47 | 0.23 | … | 1.00 |
| DDO168 | 6.03 | 0.45 | 0.00 | 4.36 | 0.76 | … | 19.71 | 0.46 | … | 2.00 |
| DDO170 | 2.45 | 0.49 | 0.00 | 3.19 | 1.84 | … | 4.92 | 0.79 | … | 2.00 |
| ESO079-G014 | 1.61 | 0.51 | 0.00 | 3.86 | 1.09 | … | 4.33 | 0.50 | … | 1.00 |
| ESO116-G012 | 1.27 | 0.55 | 0.00 | 1.03 | 1.04 | … | 2.44 | 0.35 | … | 1.00 |
| ESO444-G084 | 2.74 | 0.51 | 0.00 | 0.40 | 1.86 | … | 3.25 | 0.42 | … | 2.00 |
| ESO563-G021 | 12.37 | 0.55 | 0.00 | 16.41 | 1.00 | … | 28.84 | 0.43 | … | 1.00 |
| F561-1 | 1.14 | 0.50 | 0.00 | NAN | … | … | 1.56 | 0.52 | … | 3.00 |
| F563-1 | 0.70 | 0.50 | 0.00 | 0.95 | 2.06 | … | 1.50 | 0.56 | … | 1.00 |
| F563-V1 | 0.98 | 0.50 | 0.00 | 0.29 | 0.99 | … | 0.88 | 0.48 | … | 3.00 |
| F563-V2 | 0.47 | 0.51 | 0.00 | 0.11 | 2.20 | … | 0.99 | 0.59 | … | 1.00 |
| F565-V2 | 0.16 | 0.50 | 0.00 | 0.31 | 2.22 | … | 0.47 | 0.50 | … | 2.00 |
| F567-2 | NAN | … | … | 0.50 | 1.31 | … | 2.20 | 0.56 | … | 3.00 |
| F568-1 | 0.18 | 0.50 | 0.00 | 0.72 | 1.91 | … | 1.29 | 0.61 | … | 1.00 |
| F568-3 | 1.18 | 0.47 | 0.00 | 1.79 | 1.31 | … | 3.06 | 0.41 | … | 1.00 |
| F568-V1 | 0.19 | 0.51 | 0.00 | 0.14 | 2.14 | … | 1.04 | 0.81 | … | 1.00 |
| F571-8 | 0.75 | 0.37 | 0.00 | 2.04 | 0.17 | … | 41.61 | 0.11 | … | 1.00 |
| F571-V1 | 0.18 | 0.52 | 0.00 | 0.20 | 1.49 | … | 0.29 | 0.50 | … | 2.00 |
| F574-1 | 0.35 | 0.48 | 0.00 | 1.42 | 1.54 | … | 2.50 | 0.71 | … | 1.00 |
| F574-2 | NAN | … | … | 0.19 | 0.75 | … | 0.09 | 0.49 | … | 3.00 |
| F579-V1 | 0.06 | 0.51 | 0.00 | 1.07 | 1.63 | … | 2.56 | 0.63 | … | 1.00 |
| F583-1 | 0.36 | 0.50 | 0.00 | 1.04 | 1.87 | … | 2.66 | 0.91 | … | 1.00 |
| F583-4 | 0.42 | 0.51 | 0.00 | 0.28 | 1.30 | … | 0.13 | 0.48 | … | 1.00 |
| IC2574 | 2.51 | 0.77 | 0.00 | 2.27 | 1.10 | … | 1.44 | 0.07 | … | 2.00 |
| IC4202 | 8.67 | 0.50 | 0.31 | 32.72 | 1.09 | 0.74 | 41.91 | 1.60 | 0.34 | 1.00 |
| KK98-251 | 0.35 | 0.49 | 0.00 | 0.42 | 1.67 | … | 1.23 | 0.44 | … | 2.00 |
| NGC0024 | 0.37 | 0.54 | 0.00 | 0.73 | 1.39 | … | 0.85 | 1.01 | … | 1.00 |
| NGC0055 | 0.55 | 0.39 | 0.00 | 2.86 | 1.01 | … | 1.58 | 0.19 | … | 2.00 |
| NGC0100 | 0.10 | 0.49 | 0.00 | 0.10 | 0.93 | … | 1.29 | 0.28 | … | 1.00 |
| NGC0247 | 2.38 | 0.46 | 0.00 | 2.18 | 1.53 | … | 3.06 | 0.78 | … | 2.00 |
| NGC0289 | 1.97 | 0.45 | 0.00 | 1.78 | 0.74 | … | 2.13 | 0.92 | … | 2.00 |
| NGC0300 | 0.47 | 0.50 | 0.00 | 0.42 | 1.14 | … | 0.91 | 0.40 | … | 2.00 |
| NGC0801 | 7.07 | 0.64 | 0.00 | 7.38 | 0.77 | … | 7.75 | 1.33 | … | 1.00 |
| NGC0891 | 4.24 | 0.27 | 0.60 | 1.89 | 0.66 | 3.77E-05 | 7.37 | 0.33 | 0.40 | 1.00 |
| NGC1003 | 2.67 | 0.77 | 0.00 | 3.42 | 0.77 | … | 4.67 | 0.37 | … | 1.00 |
| NGC1090 | 1.54 | 0.53 | 0.00 | 2.27 | 0.81 | … | 2.78 | 0.74 | … | 1.00 |
| NGC1705 | 0.08 | 0.49 | 0.00 | 0.13 | 1.25 | … | 0.37 | 1.22 | … | 3.00 |
| NGC2366 | 1.09 | 0.39 | 0.00 | 2.34 | 1.06 | … | 1.93 | 0.24 | … | 3.00 |
| NGC2403 | 10.94 | 0.83 | 0.00 | 10.73 | 0.86 | … | 14.14 | 0.51 | … | 1.00 |
| NGC2683 | 1.89 | 0.60 | 0.70 | 1.01 | 0.88 | 0.44 | 3.37 | 0.55 | 0.73 | 2.00 |
| NGC2841 | 1.58 | 0.60 | 0.66 | 1.36 | 0.91 | 1.10 | 1.52 | 0.81 | 0.93 | 1.00 |
| NGC2903 | 6.67 | 0.43 | 0.00 | 7.45 | 0.62 | … | 20.64 | 0.21 | … | 1.00 |
| NGC2915 | 0.71 | 0.41 | 0.00 | 0.63 | 0.56 | … | 4.02 | 0.32 | … | 2.00 |
| NGC2955 | 4.21 | 0.36 | 0.91 | 4.45 | 0.40 | 0.88 | 3.91 | 0.37 | 0.84 | 1.00 |
| NGC2976 | 0.33 | 0.48 | 0.00 | 0.46 | 0.91 | … | 1.73 | 0.35 | … | 2.00 |
| NGC2998 | 1.47 | 0.36 | 0.00 | 3.74 | 0.87 | … | 2.94 | 0.82 | … | 1.00 |
| NGC3109 | 0.23 | 0.50 | 0.00 | 0.28 | 2.05 | … | 4.13 | 0.21 | … | 1.00 |
| NGC3198 | 1.31 | 0.52 | 0.00 | 1.72 | 0.88 | … | 2.06 | 0.77 | … | 1.00 |
| NGC3521 | 0.20 | 0.52 | 0.00 | 0.77 | 0.71 | … | 0.51 | 0.46 | … | 1.00 |
| NGC3726 | 2.54 | 0.61 | 0.00 | 2.41 | 0.76 | … | 2.98 | 0.47 | … | 2.00 |
| NGC3741 | 0.90 | 0.71 | 0.00 | 0.65 | 0.71 | … | 0.77 | 0.31 | … | 1.00 |
| NGC3769 | 0.88 | 0.43 | 0.00 | 0.71 | 0.71 | … | 0.95 | 0.41 | … | 2.00 |
| NGC3877 | 3.47 | 0.31 | 0.00 | 9.07 | 0.87 | … | 10.22 | 0.40 | … | 2.00 |
| NGC3893 | 0.69 | 0.46 | 0.00 | 0.57 | 0.73 | … | 1.00 | 0.45 | … | 1.00 |
| NGC3917 | 1.64 | 0.53 | 0.00 | 2.70 | 1.14 | … | 4.60 | 0.55 | … | 1.00 |
| NGC3949 | 0.84 | 0.52 | 0.00 | 0.76 | 0.73 | … | 0.55 | 0.44 | … | 2.00 |
| NGC3953 | 0.56 | 0.49 | 0.00 | 0.76 | 0.90 | … | 3.42 | 0.59 | … | 1.00 |
| NGC3972 | 0.80 | 0.45 | 0.00 | 2.08 | 1.03 | … | 2.07 | 0.50 | … | 1.00 |
| NGC3992 | 0.71 | 0.55 | 0.00 | 1.68 | 1.04 | … | 3.47 | 0.76 | … | 1.00 |
| NGC4010 | 2.16 | 0.46 | 0.00 | 1.90 | 0.85 | … | 2.74 | 0.36 | … | 2.00 |
| NGC4013 | 0.84 | 0.58 | 0.83 | 1.53 | 0.54 | 1.43 | 1.81 | 0.35 | 0.79 | 2.00 |
| NGC4051 | 1.73 | 0.48 | 0.00 | 1.91 | 0.81 | … | 2.49 | 0.45 | … | 2.00 |
| NGC4068 | 0.82 | 0.49 | 0.00 | 0.27 | 0.80 | … | 2.52 | 0.38 | … | 2.00 |
| NGC4085 | 5.28 | 0.36 | 0.00 | 2.83 | 0.58 | … | 9.09 | 0.35 | … | 2.00 |
| NGC4088 | 0.64 | 0.46 | 0.00 | 0.87 | 0.63 | … | 0.66 | 0.40 | … | 1.00 |
| NGC4100 | 1.21 | 0.57 | 0.00 | 1.90 | 0.95 | … | 1.66 | 0.76 | … | 1.00 |
| NGC4138 | 5.52 | 0.55 | 0.68 | 0.75 | 0.98 | 0.00 | 2.49 | 0.55 | 0.69 | 2.00 |
| NGC4157 | 0.45 | 0.53 | 0.66 | 0.61 | 0.69 | 0.56 | 0.72 | 0.43 | 0.64 | 1.00 |
| NGC4183 | 0.14 | 0.53 | 0.00 | 0.53 | 1.24 | … | 1.13 | 0.79 | … | 1.00 |
| NGC4214 | 0.77 | 0.50 | 0.00 | 1.19 | 1.01 | … | 1.06 | 0.46 | … | 2.00 |
| NGC4217 | 2.23 | 0.60 | 0.30 | 1.33 | 1.13 | 0.46 | 3.17 | 1.17 | 0.17 | 1.00 |
| NGC4389 | 6.30 | 0.29 | 0.00 | 0.20 | 0.31 | … | 9.31 | 0.30 | … | 3.00 |
| NGC4559 | 0.35 | 0.50 | 0.00 | 0.34 | 0.77 | … | 0.50 | 0.52 | … | 1.00 |
| NGC5005 | 0.10 | 0.49 | 0.60 | 0.07 | 0.62 | 0.69 | 0.09 | 0.54 | 0.56 | 1.00 |
| NGC5033 | 7.15 | 0.52 | 0.36 | 6.39 | 0.85 | 0.52 | 8.02 | 1.03 | 0.43 | 1.00 |
| NGC5055 | 6.19 | 0.26 | 0.00 | 6.29 | 0.62 | … | 7.42 | 0.56 | … | 1.00 |
| NGC5371 | 1.96 | 0.42 | 0.00 | 12.55 | 0.78 | … | 10.16 | 3.30 | … | 1.00 |
| NGC5585 | 5.18 | 0.55 | 0.00 | 6.35 | 0.75 | … | 6.82 | 0.22 | … | 1.00 |
| NGC5907 | 3.30 | 0.42 | 0.00 | 7.88 | 0.94 | … | 7.73 | 1.08 | … | 1.00 |
| NGC5985 | 2.54 | 0.53 | 0.64 | 5.91 | 1.01 | 2.12 | 6.97 | 0.63 | 3.32 | 1.00 |
| NGC6015 | 9.28 | 0.64 | 0.00 | 13.06 | 0.99 | … | 10.87 | 1.12 | … | 2.00 |
| NGC6195 | 1.81 | 0.41 | 0.79 | 2.95 | 0.41 | 0.81 | 2.26 | 0.32 | 0.85 | 1.00 |
| NGC6503 | 1.29 | 0.34 | 0.00 | 1.44 | 0.75 | … | 2.98 | 0.45 | … | 1.00 |
| NGC6674 | 1.47 | 0.87 | 0.74 | 3.98 | 0.70 | 1.87 | 10.64 | 0.95 | 1.30 | 1.00 |
| NGC6789 | NAN | … | … | 0.57 | 1.41 | … | 5.90 | 0.60 | … | 2.00 |
| NGC6946 | 1.55 | 0.57 | 0.61 | 1.75 | 0.64 | 0.58 | 1.53 | 0.64 | 0.71 | 1.00 |
| NGC7331 | 0.95 | 0.45 | 0.65 | 1.05 | 0.53 | 1.17 | 1.29 | 0.32 | 0.60 | 1.00 |
| NGC7793 | 0.79 | 0.58 | 0.00 | 0.75 | 0.89 | … | 1.01 | 0.55 | … | 1.00 |
| NGC7814 | 0.46 | 0.52 | 0.60 | 0.63 | 0.38 | 0.68 | 1.33 | 1.17 | 0.52 | 1.00 |
| PGC51017 | 5.98 | 0.41 | 0.00 | 2.31 | 0.80 | … | 4.57 | 0.44 | … | 3.00 |
| UGC00128 | 3.56 | 0.57 | 0.00 | 6.16 | 1.64 | … | 6.25 | 2.49 | … | 1.00 |
| UGC00191 | 1.97 | 0.49 | 0.00 | 2.45 | 1.38 | … | 3.84 | 1.10 | … | 1.00 |
| UGC00634 | NAN | … | … | 5.83 | 1.57 | … | 2.43 | 0.49 | … | 2.00 |
| UGC00731 | 0.22 | 0.50 | 0.00 | 0.08 | 3.79 | … | 6.42 | 2.39 | … | 1.00 |
| UGC00891 | NAN | … | … | 1.45 | 1.34 | … | 25.16 | 0.32 | … | 2.00 |
| UGC01230 | 1.18 | 0.52 | 0.00 | 0.80 | 1.71 | … | 2.95 | 0.72 | … | 1.00 |
| UGC01281 | 0.14 | 0.50 | 0.00 | 0.34 | 1.42 | … | 0.24 | 0.39 | … | 1.00 |
| UGC02023 | NAN | … | … | 0.06 | 0.72 | … | 1.15 | 0.49 | … | 2.00 |
| UGC02259 | 0.96 | 0.50 | 0.00 | 4.66 | 1.81 | … | 7.22 | 1.14 | … | 2.00 |
| UGC02455 | 1.07 | 0.45 | 0.00 | 0.90 | 0.28 | … | 6.55 | 0.33 | … | 3.00 |
| UGC02487 | 4.86 | 0.82 | 0.54 | 4.18 | 1.24 | 0.77 | 4.48 | 1.83 | 0.91 | 1.00 |
| UGC02885 | 0.84 | 0.47 | 1.00 | 2.14 | 0.50 | 0.98 | 0.86 | 0.45 | 0.97 | 1.00 |
| UGC02916 | 11.58 | 1.06 | 0.50 | 11.36 | 1.41 | 0.71 | 11.65 | 1.57 | 0.73 | 2.00 |
| UGC02953 | 4.97 | 0.50 | 0.60 | 5.69 | 0.74 | 0.78 | 5.66 | 0.61 | 0.62 | 2.00 |
| UGC03205 | 3.82 | 0.67 | 0.88 | 2.99 | 0.68 | 1.07 | 4.20 | 0.73 | 1.32 | 1.00 |
| UGC03546 | 0.99 | 0.61 | 0.53 | 1.14 | 0.61 | 0.60 | 0.91 | 0.68 | 0.51 | 1.00 |
| UGC03580 | 2.40 | 0.94 | 0.32 | 2.20 | 0.48 | 0.41 | 2.29 | 0.29 | 0.11 | 2.00 |
| UGC04278 | 0.58 | 0.66 | 0.00 | 0.92 | 1.00 | … | 2.60 | 0.53 | … | 1.00 |
| UGC04305 | 1.45 | 0.54 | 0.00 | NAN | … | … | 2.02 | 0.71 | … | 3.00 |
| UGC04325 | 2.89 | 0.52 | 0.00 | 3.72 | 1.87 | … | 9.43 | 0.94 | … | 1.00 |
| UGC04483 | 0.54 | 0.50 | 0.00 | 0.52 | 1.28 | … | 0.87 | 0.43 | … | 2.00 |
| UGC04499 | 0.46 | 0.46 | 0.00 | 1.54 | 1.14 | … | 1.78 | 0.51 | … | 1.00 |
| UGC05005 | 0.05 | 0.50 | 0.00 | 0.09 | 1.03 | … | 0.32 | 0.45 | … | 1.00 |
| UGC05253 | 3.79 | 0.55 | 0.72 | 5.65 | 0.53 | 0.72 | 4.75 | 0.63 | 0.69 | 2.00 |
| UGC05414 | 0.37 | 0.52 | 0.00 | 0.19 | 1.06 | … | 1.30 | 0.41 | … | 1.00 |
| UGC05716 | 3.07 | 0.50 | 0.00 | 3.81 | 1.52 | … | 5.66 | 1.41 | … | 2.00 |
| UGC05721 | 0.94 | 0.50 | 0.00 | 0.84 | 1.11 | … | 1.82 | 0.62 | … | 1.00 |
| UGC05750 | 0.26 | 0.49 | 0.00 | 0.47 | 1.57 | … | 1.35 | 0.48 | … | 1.00 |
| UGC05764 | 7.05 | 0.49 | 0.00 | 9.45 | 3.70 | … | 16.18 | 3.83 | … | 2.00 |
| UGC05829 | 0.19 | 0.51 | 0.00 | 0.07 | 1.85 | … | 0.45 | 0.60 | … | 2.00 |
| UGC05918 | 0.03 | 0.50 | 0.00 | 0.21 | 2.31 | … | 0.94 | 0.54 | … | 2.00 |
| UGC05986 | 1.72 | 0.56 | 0.00 | 1.69 | 1.15 | … | 4.00 | 0.31 | … | 2.00 |
| UGC05999 | NAN | … | … | 3.56 | 1.27 | … | 5.69 | 0.48 | … | 2.00 |
| UGC06399 | 0.15 | 0.51 | 0.00 | 0.22 | 1.43 | … | 0.52 | 0.53 | … | 1.00 |
| UGC06446 | 0.22 | 0.50 | 0.00 | 0.17 | 1.69 | … | 1.00 | 1.04 | … | 1.00 |
| UGC06614 | 0.18 | 0.50 | 0.71 | 1.20 | 0.72 | 0.59 | 1.16 | 0.51 | 0.50 | 1.00 |
| UGC06628 | 0.31 | 0.50 | 0.00 | NAN | … | … | 0.85 | 0.52 | … | 2.00 |
| UGC06667 | 0.39 | 0.51 | 0.00 | 2.41 | 3.77 | … | 5.36 | 1.00 | … | 1.00 |
| UGC06786 | 0.69 | 0.35 | 0.78 | 0.69 | 0.50 | 0.66 | 1.39 | 0.27 | 0.34 | 1.00 |
| UGC06787 | 18.09 | 0.90 | 0.19 | 24.78 | 0.73 | 0.57 | 20.81 | 0.45 | 0.28 | 2.00 |
| UGC06818 | 2.10 | 0.45 | 0.00 | 1.24 | 0.53 | … | 5.39 | 0.29 | … | 2.00 |
| UGC06917 | 0.39 | 0.48 | 0.00 | 1.09 | 1.12 | … | 1.32 | 0.54 | … | 1.00 |
| UGC06923 | 1.97 | 0.48 | 0.00 | 0.88 | 0.80 | … | 1.62 | 0.42 | … | 2.00 |
| UGC06930 | 0.22 | 0.52 | 0.00 | 0.55 | 1.25 | … | 1.23 | 0.63 | … | 1.00 |
| UGC06973 | 1.45 | 0.31 | 0.60 | 0.41 | 0.36 | 0.81 | 15.58 | 0.17 | 0.39 | 3.00 |
| UGC06983 | 0.58 | 0.51 | 0.00 | 0.77 | 1.26 | … | 1.39 | 0.77 | … | 1.00 |
| UGC07089 | 0.23 | 0.51 | 0.00 | 0.15 | 0.95 | … | 0.43 | 0.36 | … | 2.00 |
| UGC07125 | 0.40 | 0.50 | 0.00 | 1.18 | 1.08 | … | 1.60 | 0.92 | … | 1.00 |
| UGC07151 | 1.58 | 0.56 | 0.00 | 1.30 | 1.15 | … | 3.75 | 0.50 | … | 1.00 |
| UGC07232 | NAN | … | … | 0.76 | 0.82 | … | 6.17 | 0.46 | … | 2.00 |
| UGC07261 | 0.16 | 0.47 | 0.00 | 1.46 | 1.22 | … | 0.83 | 0.56 | … | 2.00 |
| UGC07323 | 0.43 | 0.51 | 0.00 | 0.26 | 0.97 | … | 0.66 | 0.41 | … | 1.00 |
| UGC07399 | 0.36 | 0.50 | 0.00 | 1.00 | 1.54 | … | 1.90 | 0.59 | … | 1.00 |
| UGC07524 | 0.34 | 0.53 | 0.00 | 1.48 | 1.59 | … | 1.84 | 0.79 | … | 1.00 |
| UGC07559 | 0.26 | 0.50 | 0.00 | 0.21 | 0.97 | … | 2.60 | 0.31 | … | 2.00 |
| UGC07577 | 0.22 | 0.49 | 0.00 | 0.06 | 0.67 | … | 5.79 | 0.24 | … | 2.00 |
| UGC07603 | 0.47 | 0.48 | 0.00 | 0.52 | 1.07 | … | 1.77 | 0.34 | … | 1.00 |
| UGC07608 | 0.14 | 0.50 | 0.00 | 0.30 | 2.02 | … | 0.73 | 0.48 | … | 1.00 |
| UGC07690 | 0.87 | 0.53 | 0.00 | 0.40 | 1.05 | … | 1.53 | 0.60 | … | 2.00 |
| UGC07866 | 0.12 | 0.50 | 0.00 | 0.07 | 1.26 | … | 0.26 | 0.45 | … | 2.00 |
| UGC08286 | 1.00 | 0.52 | 0.00 | 2.66 | 1.66 | … | 2.64 | 1.05 | … | 1.00 |
| UGC08490 | 0.21 | 0.51 | 0.00 | 0.15 | 1.26 | … | 0.34 | 0.86 | … | 1.00 |
| UGC08550 | 0.55 | 0.50 | 0.00 | 0.70 | 1.35 | … | 1.55 | 0.74 | … | 1.00 |
| UGC08699 | 0.87 | 0.75 | 0.59 | 0.95 | 0.55 | 0.76 | 0.99 | 0.63 | 0.70 | 2.00 |
| UGC08837 | 0.68 | 0.49 | 0.00 | 0.69 | 0.78 | … | 2.35 | 0.20 | … | 2.00 |
| UGC09037 | 1.09 | 0.30 | 0.00 | 1.55 | 0.58 | … | 2.26 | 0.20 | … | 2.00 |
| UGC09133 | 6.78 | 0.65 | 0.53 | 7.50 | 0.77 | 0.72 | 6.94 | 1.64 | 1.10 | 1.00 |
| UGC09992 | NAN | … | … | 0.04 | 1.30 | … | 1.08 | 0.51 | … | 2.00 |
| UGC10310 | 0.64 | 0.51 | 0.00 | 0.12 | 1.54 | … | 1.76 | 0.62 | … | 1.00 |
| UGC11455 | 2.40 | 0.41 | 0.00 | 3.98 | 0.76 | … | 6.55 | 0.38 | … | 1.00 |
| UGC11557 | 0.67 | 0.48 | 0.00 | 0.91 | 0.60 | … | 3.18 | 0.42 | … | 2.00 |
| UGC11820 | 3.46 | 1.19 | 0.00 | 1.44 | 1.24 | … | 1.99 | 1.01 | … | 1.00 |
| UGC11914 | 0.65 | 0.42 | 0.75 | 1.61 | 0.07 | 0.89 | 1.73 | 0.22 | 0.48 | 1.00 |
| UGC12506 | 0.29 | 0.59 | 0.00 | 1.18 | 1.35 | … | 1.98 | 1.12 | … | 2.00 |
| UGC12632 | 0.17 | 0.50 | 0.00 | 0.21 | 1.86 | … | 1.80 | 1.08 | … | 1.00 |
| UGC12732 | 0.24 | 0.48 | 0.00 | 0.18 | 1.51 | … | 0.50 | 1.07 | … | 1.00 |
| UGCA281 | 0.38 | 0.49 | 0.00 | 0.28 | 1.23 | … | 0.47 | 0.37 | … | 3.00 |
| UGCA442 | 1.60 | 0.50 | 0.00 | 0.86 | 2.28 | … | 7.65 | 0.44 | … | 1.00 |
| UGCA444 | 0.22 | 0.50 | 0.00 | 0.10 | 3.78 | … | 0.33 | 0.49 | … | 2.00 |
Data Availability
The code used to generate all figures and fit results is publicly available at https://github.com/Cisneros-Galaxy/RCFM. The rotation curve data and photometric models used in this paper are from the SPARC database available at http://astroweb.cwru.edu/SPARC/. The few other datasets used as examples are available at the cited reference.